# Understanding the differences in Foundation Models: Attention, State Space Models, and Recurrent Neural Networks

### v2

@@ 6 Conclusion
In this paper we presented the DSF, a framework based on dynamical-systems theory that allows analysis of different deep learning architectures by writing them as linear recurrences in state space. We first showed how to reformulate different architectures into the DSF, and then explored (theoretical and computational) insights resulting from this analysis. For instance, we showed that with proper normalization the performance of linear attention can be significantly increased (see Fig. 2). We also show, that the DSF allows to integrate insights from one architecture to another as exemplified by Section 4.2. Lastly, the DSF can provide a deeper theoretical understanding on the relationship between softmax attention and other models. We expect that the DSF can serve as a tool for principled analysis and design of deep learning architectures.
@@
In terms of the limitations of the DSF, it is important to highlight that, while the DSF parametrization allows for a principled comparison between frameworks, architectures in DSF do not necessarily enjoy an efficient implementation unless their specific structure can leverage some of the existing algorithms (parallel scan, etc.). In terms of the computational experiments, the insights mentioned above are only verified on a synthetic task (MQAR) and for some on a small size language task (WikiText-103). To strengthen the insights, a more in-depth analysis is needed on a larger and more complex language task.
Acknowledgements
@@ Appendix A Visual Representation of the matrix dimensions
where $h_{i}\in\mathbb{R}^{N}$ is the hidden state, $\Lambda_{i}\in\mathbb{R}^{N\times N}$ is the diagonal state transition matrix, $B_{i}\in\mathbb{R}^{N\times d}$ is the input matrix. We highlight the role of the state expansion, where $u\in\mathbb{R}^{d}$ and $h\in\mathbb{R}^{N}=\mathbb{R}^{nd}$.
@@ Appendix B Derivation of Separable Attention into DSF
$y_{i}=\sum_{j=0}^{i}f(q_{i},k_{j},\mathbf{k}_{i})W_{V}u_{j}$
where we define the sequence of keys
@@
We show that it can be equivalently written as the LTV system (11), i.e.,
@@
with $\Lambda_{i}\in\mathbb{R}^{nd\times nd}$, $B_{i}\in\mathbb{R}^{nd\times d}$ and $C_{i}\in\mathbb{R}^{d\times nd}$, if the function $f(\cdot,\cdot)$ can be separated as follows
@@
where $\phi(q_{i})\in\mathbb{R}^{n}$, $\psi(k_{j})\in\mathbb{R}^{n}$, and $\eta(q_{i},\mathbf{k}_{i})\in\mathbb{R}$ can be considered to be a normalization. If the unnormalized part of $f(\cdot,\cdot)$ is a kernel, it holds that $\psi(k_{j})=\phi(k_{j})$ for some, possibly infinite dimensional, feature vector $\phi$.
@@
$\Rightarrow$ $B_{2}=\left(\frac{1}{\eta(q_{2},\mathbf{k}_{2})}\mathbb{I}_{d}\otimes\psi(k_{2})\right)W_{V},\;C_{2}=\mathbb{I}_{d}\otimes\phi(q_{2})^{\top}\textrm{ and }\Lambda_{2}=\frac{\eta(q_{1},\mathbf{k}_{1})}{\eta(q_{2},\mathbf{k}_{2})}\mathbb{I}_{nd}$
Plugging it back in we can see
@@
$B_{i}$ $=\left(\frac{1}{\eta(q_{i},\mathbf{k}_{i})}\mathbb{I}_{d}\otimes\psi(k_{i})\right)W_{V}$
$C_{i}$ $=\mathbb{I}_{d}\otimes\phi(q_{i})^{\top}$
$\Lambda_{i}$ $=\frac{\eta(q_{i-1},\mathbf{k}_{i-1})}{\eta(q_{i},\mathbf{k}_{i})}\mathbb{I}_{nd}$
@@ Appendix C Proof of Lemma 1: Derivation of Softmax Attention into DSF
The function $\textup{softmax}:\mathbb{R}^{n\times m}\rightarrow(0,1]^{n\times m}$ is defined through row-wise normalisation and is given by
@@
where $\textup{softmax}_{i,j}$ refers to the element $(i,j)$ of the corresponding matrix. Note that $e^{q_{i}^{\top}k_{j}}$ is a kernel, implying that the softmax attention can be brought into the separable form (13). In order to provide the separable form for $e^{q_{i}^{\top}k_{j}}$, we first consider the Taylor expansion of the kernel, which is given by
@@
Each polynomial $(q_{i}^{\top}k_{j})^{p}$ represents itself a homogeneous polynomial kernel and its decomposition into a feature vector of $\left(\begin{matrix}n+p-1\\ p\end{matrix}\right)$ monomials, as shown in [Shashua, 2009], is given by
@@
$e^{q_{i}^{\top}k_{j}}$ $=\begin{bmatrix}1,q_{i},\sqrt{\frac{1}{2!}}\tilde{\phi}_{2}(q_{i})^{\top},\sqrt{\frac{1}{3!}}\tilde{\phi}_{3}(q_{i})^{\top},\dots\end{bmatrix}\begin{bmatrix}1,k_{j}^{\top},\sqrt{\frac{1}{2!}}\tilde{\phi}_{2}(k_{j})^{\top},\sqrt{\frac{1}{3!}}\tilde{\phi}_{3}(k_{j})^{\top},\dots\end{bmatrix}^{\top}$ (24)
@@
Note that to attain the monomials in (23) for a given $p$, one can also use $\bigotimes_{j=1}^{p}x$, as given in, e.g., [López-Sánchez et al., 2018], which is equivalent up to the constant coefficients, such that we can use $\sqrt{\frac{1}{p!}}\tilde{\phi}_{p}(x)=c_{p}\bigotimes_{j=1}^{p}x$ where $c_{p}$ is a matrix of the respective coefficients multiplying each monomial.
Appendix D Derivation of S6 into DSF
While $A$ in (4) is represented as a dense matrix of size $n\times d$ purely for computational reasons, mathematically $A$ is a diagonal matrix of size $nd\times nd$. This is evident from the fact that S6 parameterizes a different submatrix $A^{d}\in\mathbb{R}^{n\times n}$ for each embedding dimension, leading to
@@ Appendix D Derivation of S6 into DSF
$\Delta_{i}=\textrm{diag}(\sigma(W_{\Delta}(W_{u}u_{i})+b_{\Delta})),$ (25)
where $\sigma(\cdot)$ denotes the softplus function, $W_{u}\in\mathbb{R}^{p\times d},\,W_{\Delta}\in\mathbb{R}^{d\times p}$ are weight matrices with $p<d$, and $b_{\Delta}\in\mathbb{R}^{d}$ is a bias. Note that we embed the computed discretization times in a diagonal $d\times d$ matrix to simplify the next reformulations. The product of $\Delta_{i}$ and $A$ is performed along the embedding dimension axis, i.e.,
@@
Appendix E Derivation of LSTMs into DSF
E.1 RG-LRU
@@
As with S6 in Appendix D, we reformulate some quantities for easier presentation, namely we embed the input-dependent vectors $\sigma(W_{R}u_{i})\in\mathbb{R}^{d},\,\sigma(W_{B}u_{i})\in\mathbb{R}^{d}$, where $\sigma(\cdot)$ denotes the sigmoid function, in a diagonal matrix, i.e.,
@@
Appendix F Dynamical System Derivation of Multi-headed Separable Attention
@@
Additionally, we consider the multi-headed setting introduced in [Vaswani et al., 2017, Section 3.2.2], i.e., $s$ different attention operations are performed in parallel. Due to the right multiplication of the input (instead of left multiplication as in the original paper), the output of the different heads is stacked row-wise instead of column-wise. Additionally, we assume there is no output mapping after the attention operation. This yields the simplified multi-headed layer
@@
$=\sum_{j=0}^{i}\begin{bmatrix}[\phi(q_{i})^{\top}]^{1}&&\\ &\ddots&\\ &&[\phi(q_{i})^{\top}]^{s}\end{bmatrix}\begin{bmatrix}[\frac{1}{\eta(q_{i},\mathbf{k}_{i})}]^{1}\cdot\mathbb{I}_{n/h}&&\\ &\ddots&\\ &&[\frac{1}{\eta(q_{i},\mathbf{k}_{i})}]^{s}\cdot\mathbb{I}_{n/h}\end{bmatrix}$
@@
where $\cdot^{r}$ denotes the head index. As is standard for multi-headed attention, we reduce the dimensions of the queries, keys, and values by the number of heads, i.e., $q_{i}\in\mathbb{R}^{m/h}$, $k_{j}\in\mathbb{R}^{m/h}$, and $v_{j}\in\mathbb{R}^{d/h}$. Since the $h$ different values $[v_{j}]^{s}$ are stacked, this is equivalent to the single headed version, i.e.,
@@
Above observation is also valid for the queries and keys, i.e., we can e.g. write $[\phi(q_{i})]^{s}$ using an indicator function $\mathcal{I}(\cdot)$ on the single headed query $q_{i}$, i.e.,
$[\phi(q_{i})]^{s}=\phi(\mathcal{I}_{h}(q_{i}))=\phi(\mathcal{I}_{h}(W_{Q}u_{i})).$
@@
$\Lambda_{i}$ $=\textrm{diag}\left(\frac{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{1}}{[\eta(q_{i},\mathbf{k}_{i})]^{1}}\mathbb{I}_{d/h},\dots,\frac{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{s}}{[\eta(q_{i},\mathbf{k}_{i})]^{s}}\mathbb{I}_{d/h}\right)\otimes\mathbb{I}_{n}\in\mathbb{R}^{nd\times nd},$ (27a)
$B_{i}$ $=\left[\textrm{diag}\left(\frac{1}{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{1}}\mathbb{I}_{d/h},\dots,\frac{1}{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{s}}\mathbb{I}_{d/h}\right)\otimes\psi(\mathcal{I}_{h}(k_{i}))\right]W_{V}\in\mathbb{R}^{nd\times d},$ (27b)
$C_{i}$ $=\mathbb{I}_{d}\otimes\phi(\mathcal{I}_{h}(q_{i}))^{\top}\in\mathbb{R}^{d\times nd}.$ (27c)
@@
Appendix G S6 uses reversed sigmoid in state transition matrix
@@
Appendix H Proof of Lemma 2
$\Lambda^{\bar{N}}_{i}=\begin{bmatrix}\Lambda^{N}_{i}&0\\ \tilde{\Lambda}&\bar{\Lambda}\end{bmatrix},\;B^{\bar{N}}_{i}=\begin{bmatrix}B^{N}_{i}\\ \tilde{B}\end{bmatrix},$
$C^{\bar{N}}_{i}=\begin{bmatrix}C^{N}_{i}&0\end{bmatrix},\;D^{\bar{N}}_{i}=D^{N}_{i}.$
It can be seen that the $N$ first states of the system with dimension $\bar{N}$ propagate equivalently to the states of the system of dimension $N$. The additional states can evolve independently given any matrices $\tilde{\Lambda}$, $\bar{\Lambda}$ and $\tilde{B}$ and do not affect the output, such that both systems are equivalent. The two outputs are then equivalent by setting the corresponding entries of the $C^{\bar{N}}_{i}$ matrix to 00.
Appendix I Experimental Details
The experimental results provided in Section 4 are performed on the multi-query associative recall (MQAR) benchmark proposed in [Arora et al., 2023] and the WikiText-103 dataset. To obtain the MQAR results, we modified the Zoology code base [^5] https://github.com/HazyResearch/zoology [Arora et al., 2023] and added the missing models. We provide our code here: https://github.com/IntelligentControlSystems/dsf-mqar.
I.1 MQAR experiments
@@
    Attention: softmax attention [Vaswani et al., 2017], linear attention [Katharopoulos et al., 2020], normalized attention (21). For all three attention functions, we use a standard GPT-2 style multi-headed Transformer architecture, where we replace the attention block with the respective attention function. The three attention functions are defined in Section 2.1. For all MQAR runs we use a single attention head.
@@
    State space model: S6 [Gu and Dao, 2023], SSD [Dao and Gu, 2024]. For both SSM variants, we use a standard GPT-2 style single-headed Transformer architecture, where we replace the attention block with the respective SSM variant. This means for S6 and SSD, we do not implement the pre-convolution on the input or the SiLU activations; but just the S6 and SSD blocks. We do this to ensure a fair comparison of the backbones (sequence mixers) irrespective of the higher-level architecture. The S6 block is defined in Section 2.2 and we use the provided code base [^6] https://github.com/state-spaces/mamba to implement it. For SSD there is no exact mathematical definition in the paper [Dao and Gu, 2024] and no code provided. Therefore, we modify the S6 architecture and code to incorporate the ideas presented in [Dao and Gu, 2024]. This means we replace the learnt parameter $A\in\mathbb{R}^{n\times d}$ with a single learnt scalar $a$ and reduce the dimension of the input-dependent discretization time $\Delta_{i}$ from $d$ to a scalar as well. We then implement this with the selective scan algorithm [Gu and Dao, 2023] instead of the newly proposed SSD algorithm.
@@
    RNN: qLSTM [Stanić et al., 2023], modified qLSTM. For implementation of the qLSTM, we build on the Hippogriff codebase. [^7] https://github.com/proger/hippogriff We embed both qLSTM variants in a standard GPT-2 style single-headed Transformer architecture, where we replace the attention block with the qLSTM. We do this to ensure a fair comparison of the backbones (sequence mixers) irrespective of the higher-level architecture. The standard qLSTM is defined in Section 2.3 and the modified qLSTM is the same as the standard qLSTM but with modified state transition (22), i.e., a modified forget gate.
@@
  • Width and depth: For all runs, we use two layers (each with a sequence model and a MLP, interleaved with layer normalization). The model dimensions $d$, state expansion $n$, sequence length $L$, and number of KV pairs are swept according to the experiment (see Section 4). This is the same setup as in [Arora et al., 2023].
  • Position information: Positional embeddings [Brown et al., 2020] are used for the attention and RNN architecture classes, but not for the SSM architecture classes. This is the same setup as in [Arora et al., 2023].
  • Data: Each model is trained on 100,000 datapoints and evaluated on 3,000 datapoints. The data and its order are constant for all runs. This is the same setup as in [Arora et al., 2023].
@@ I.2 WikiText-103 experiments
We use the 70M parameter Pythia architecture (Pythia70M) [Biderman et al., 2023]. [^9] https://github.com/EleutherAI/pythia For softmax attention we use the standard Pythia attention block, while for linear attention [Katharopoulos et al., 2020] and normalized attention (21) we replace the attention block in the Pythia architecture with the respective attention functions defined in Sections 2.1 and 4.2.

### v3

@@ 6 Conclusion
In this paper we presented the DSF, a framework based on dynamical-systems theory that allows analysis of different deep learning architectures by writing them as linear recurrences in state space. We first showed how to reformulate different architectures into the DSF, and then explored (theoretical and experimental) insights resulting from this analysis, thereby answering the questions posed in the introduction. For instance, we showed that with proper normalization the performance of linear attention can be significantly increased (see Fig. 2). We also show, that the DSF allows to integrate insights from one architecture to another as exemplified by Section 4.2. Additionally, the DSF naturally allows analysis of the eigenvalues of the state transition matrix $A$, which are linked to the exploding/vanishing gradient problem [6]. In the case of SSMs and RNNs, the eigenvalues are constrained to be stable by construction, for attention-based models this is not the case and stability needs to be obtained via normalization. The eigenvalues together with the state expansion also affect a model’s long-term memory [6]. Both of these aspects can be analyzed via the DSF and should be further investigated in future work. While the training dynamics (especially convergence) can be studied empirically using experiments, the DSF also allows a theoretical analysis. As discussed in Example 2 of [56], a gradient-based optimization algorithm (e.g. SGD) can be interpreted and written as a dynamical system. Using this viewpoint together with the DSF allows interpretation of the training dynamics as two interacting dynamical systems. Therefore, the training dynamics can be theoretically analyzed using tools from control theory, e.g., via Lyapunov theory for convergence and stability of the training. However, we believe this question requires an in-depth investigation and additional empirical validation of the theoretical findings. We expect that the DSF can serve as a tool for principled analysis and design of deep learning architectures.
@@
In terms of limitations, it is important to highlight that, while the DSF parametrization allows for a principled comparison between frameworks, architectures written in the DSF do not necessarily enjoy an efficient implementation unless their specific structure can leverage some of the existing algorithms (parallel scan, etc.). In terms of experiments, the insights mentioned above are only verified on two synthetic tasks (MQAR/LRA) and a smaller language task (WikiText-103). To strengthen the insights, a more detailed analysis is needed on larger and more complex tasks.
Acknowledgments and Disclosure of Funding
@@ Appendix A Visual Representation of the matrix dimensions
where $h_{i}\in\mathbb{R}^{N}$ is the hidden state, $\Lambda_{i}\in\mathbb{R}^{N\times N}$ is the diagonal state transition matrix, and $B_{i}\in\mathbb{R}^{N\times d}$ is the input matrix. We highlight the role of the state expansion, where $u\in\mathbb{R}^{d}$ and $h\in\mathbb{R}^{N}=\mathbb{R}^{nd}$.
@@ Appendix B Derivation of Separable Attention into DSF
$y_{i}=\sum_{j=0}^{i}f(q_{i},k_{j},\mathbf{k}_{i})W_{V}u_{j},$
where we define the sequence of keys as
@@
We show that this layer can be equivalently written as the LTV system (11), i.e.,
@@
with $\Lambda_{i}\in\mathbb{R}^{nd\times nd}$, $B_{i}\in\mathbb{R}^{nd\times d}$ and $C_{i}\in\mathbb{R}^{d\times nd}$, if the function $f(\cdot,\cdot)$ is separable as follows
@@
where $\phi(q_{i})\in\mathbb{R}^{n}$, $\psi(k_{j})\in\mathbb{R}^{n}$, and $\eta(q_{i},\mathbf{k}_{i})\in\mathbb{R}$ can be considered to be a normalization function. If the unnormalized part of $f(\cdot,\cdot)$ is a kernel function, it holds that $\psi(k_{j})=\phi(k_{j})$ for some, possibly infinite dimensional, feature vector $\phi$.
@@
$\Rightarrow$ $B_{2}=\left(\frac{1}{\eta(q_{2},\mathbf{k}_{2})}\mathbb{I}_{d}\otimes\psi(k_{2})\right)W_{V},\;C_{2}=\mathbb{I}_{d}\otimes\phi(q_{2})^{\top}\textrm{ and }\Lambda_{2}=\frac{\eta(q_{1},\mathbf{k}_{1})}{\eta(q_{2},\mathbf{k}_{2})}\mathbb{I}_{nd}.$
Plugging this back in, we observe
@@
$B_{i}$ $=\left(\frac{1}{\eta(q_{i},\mathbf{k}_{i})}\mathbb{I}_{d}\otimes\psi(k_{i})\right)W_{V},$
$C_{i}$ $=\mathbb{I}_{d}\otimes\phi(q_{i})^{\top},$
$\Lambda_{i}$ $=\frac{\eta(q_{i-1},\mathbf{k}_{i-1})}{\eta(q_{i},\mathbf{k}_{i})}\mathbb{I}_{nd}.$
@@ Appendix C Proof of Lemma 1: Derivation of Softmax Attention into DSF
The softmax function $\mathbb{R}^{n\times m}\rightarrow(0,1]^{n\times m}$ is defined through row-wise normalization and is given by
@@
where $\textup{softmax}_{i,j}(\cdot)$ refers to the element $(i,j)$ of the corresponding matrix. Note that $e^{q_{i}^{\top}k_{j}}$ is a kernel, implying that softmax attention can be brought into the separable form (13). In order to provide the separable form of $e^{q_{i}^{\top}k_{j}}$, we first consider the Taylor expansion of the kernel, which is given by
@@
Each polynomial $(q_{i}^{\top}k_{j})^{p}$ represents itself a homogeneous polynomial kernel and its decomposition into a feature vector of $\left(\begin{matrix}n+p-1\\ p\end{matrix}\right)$ monomials, as shown in [57], is given by
@@
$e^{q_{i}^{\top}k_{j}}$ $=\begin{bmatrix}1,q_{i},\sqrt{\frac{1}{2!}}\tilde{\phi}_{2}(q_{i})^{\top},\sqrt{\frac{1}{3!}}\tilde{\phi}_{3}(q_{i})^{\top},\dots\end{bmatrix}\begin{bmatrix}1,k_{j},\sqrt{\frac{1}{2!}}\tilde{\phi}_{2}(k_{j})^{\top},\sqrt{\frac{1}{3!}}\tilde{\phi}_{3}(k_{j})^{\top},\dots\end{bmatrix}^{\top}$ (25)
@@
Note that to attain the monomials in (24) for a given $p$, one can also use $\bigotimes_{j=1}^{p}x$, as given in, e.g., [58], which is equivalent up to the constant coefficients, such that we can use $\sqrt{\frac{1}{p!}}\tilde{\phi}_{p}(x)=c_{p}\bigotimes_{j=1}^{p}x$, where $c_{p}$ is a matrix of the respective coefficients multiplying each monomial.
Appendix D Proof of Lemma 2
$\Lambda^{\bar{N}}_{i}=\begin{bmatrix}\Lambda^{N}_{i}&0\\ \bar{\Lambda}&\hat{\Lambda}\end{bmatrix},\;B^{\bar{N}}_{i}=\begin{bmatrix}B^{N}_{i}\\ \bar{B}\end{bmatrix},$
$C^{\bar{N}}_{i}=\begin{bmatrix}C^{N}_{i}&0\end{bmatrix},\;D^{\bar{N}}_{i}=D^{N}_{i}.$
It can be seen that the $N$ first states of the system with dimension $\bar{N}$ propagate equivalently to the states of the system of dimension $N$. The additional states evolve independently given any matrices $\bar{\Lambda}$, $\hat{\Lambda}$, $\bar{B}$ of appropriate dimension and do not affect the output, such that both systems are equivalent. The two outputs are then equivalent by setting the corresponding entries of the $C^{\bar{N}}_{i}$ matrix to 00.
Appendix E Derivation of S6 into DSF
While $A$ in (4) is represented as a dense matrix of size $n\times d$ purely for computational reasons, mathematically $A$ is a diagonal matrix of size $nd\times nd$. This is evident from the fact that S6 parameterizes a different submatrix $A^{d}\in\mathbb{R}^{n\times n}$ for each embedding dimension $d$, leading to
@@ Appendix E Derivation of S6 into DSF
$\Delta_{i}=\textrm{diag}(\textrm{softplus}(W_{\Delta}(W_{u}u_{i})+b_{\Delta})),$ (26)
where $W_{u}\in\mathbb{R}^{p\times d},\,W_{\Delta}\in\mathbb{R}^{d\times p}$ are weight matrices with $p<d$, and $b_{\Delta}\in\mathbb{R}^{d}$ is a bias. Note that we embed the computed discretization times in a diagonal $d\times d$ matrix to simplify the next reformulations. The product of $\Delta_{i}$ and $A$ is performed along the embedding dimension axis, i.e.,
@@
Appendix F Derivation of LSTMs into DSF
F.1 RG-LRU
@@
As with S6 in Appendix E, we reformulate some quantities for easier presentation, namely we embed the input-dependent vectors $\sigma(W_{R}u_{i})\in\mathbb{R}^{d},\,\sigma(W_{B}u_{i})\in\mathbb{R}^{d}$, where $\sigma(\cdot)$ denotes the sigmoid function, in a diagonal matrix, i.e.,
@@
Appendix G Dynamical System Derivation of Multi-headed Separable Attention
@@
Additionally, we consider the multi-headed setting introduced in [2, Section 3.2.2], i.e., $s$ different attention operations are performed in parallel. Due to the right multiplication of the input (instead of left multiplication as in the original paper), the output of the different heads is stacked row-wise instead of column-wise. Additionally, we assume there is no output mapping after the attention operation. This yields the simplified multi-headed layer
@@
$=\sum_{j=0}^{i}\begin{bmatrix}[\phi(q_{i})^{\top}]^{1}&&\\ &\ddots&\\ &&[\phi(q_{i})^{\top}]^{s}\end{bmatrix}\begin{bmatrix}[\frac{1}{\eta(q_{i},\mathbf{k}_{i})}]^{1}\cdot\mathbb{I}_{n/s}&&\\ &\ddots&\\ &&[\frac{1}{\eta(q_{i},\mathbf{k}_{i})}]^{s}\cdot\mathbb{I}_{n/s}\end{bmatrix}$
@@
where $\cdot^{s}$ denotes the head index. As is standard for multi-headed attention, we reduce the dimensions of the queries, keys, and values by the number of heads, i.e., $q_{i}\in\mathbb{R}^{m/s}$, $k_{j}\in\mathbb{R}^{m/s}$, and $v_{j}\in\mathbb{R}^{d/s}$. Since the $s$ different values $[v_{j}]^{s}$ are stacked, this is equivalent to the single headed version, i.e.,
@@
Above observation is also valid for the queries and keys, i.e., we can e.g. write $[\phi(q_{i})]^{s}$ using an indicator function $\mathcal{I}_{s}(\cdot)$ on the single headed query $q_{i}$, i.e.,
$[\phi(q_{i})]^{s}=\phi(\mathcal{I}_{s}(q_{i}))=\phi(\mathcal{I}_{s}(W_{Q}u_{i})).$
@@
$\Lambda_{i}$ $=\textrm{diag}\left(\frac{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{1}}{[\eta(q_{i},\mathbf{k}_{i})]^{1}}\mathbb{I}_{d/s},\dots,\frac{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{s}}{[\eta(q_{i},\mathbf{k}_{i})]^{s}}\mathbb{I}_{d/s}\right)\otimes\mathbb{I}_{n}\in\mathbb{R}^{nd\times nd},$ (28a)
$B_{i}$ $=\left[\textrm{diag}\left(\frac{1}{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{1}}\mathbb{I}_{d/s},\dots,\frac{1}{[\eta(q_{i-1},\mathbf{k}_{i-1})]^{s}}\mathbb{I}_{d/s}\right)\otimes\psi(\mathcal{I}_{s}(k_{i}))\right]W_{V}\in\mathbb{R}^{nd\times d},$ (28b)
$C_{i}$ $=\mathbb{I}_{d}\otimes\phi(\mathcal{I}_{s}(q_{i}))^{\top}\in\mathbb{R}^{d\times nd}.$ (28c)
@@
Appendix H Alternative Normalization Schemes
For all experiments in Section 4.2, we use the normalization scheme in (22). The exponential normalization function $\eta(u_{i})$ is inspired by softmax attention and S6, which both use exponential functions for normalization (see (18) and (19)). However, other normalization functions can also be considered e.g.
$\eta(u_{i})$ $=\textrm{softplus}(W_{\eta}u_{i}),$ (29)
$\eta(u_{i})$ $=\sigma(W_{\eta}u_{i}),$ (30)
[…]
Normalization Function Task [%]
[…]
Exponential (22) 35.96 85.9
Softplus (29) 35.27 84.3
Sigmoid (30) 35.80 84.7
Appendix I S6 uses reversed sigmoid in state transition matrix
@@
Appendix J Experimental Details
The experimental results provided in Section 4 are performed on the multi-query associative recall (MQAR) benchmark [24], the long range arena (LRA) benchmark [3], and the WikiText-103 dataset. To obtain the MQAR and LRA results, we modified the Zoology [^5] https://github.com/HazyResearch/zoology and LRA [^6] https://github.com/google-research/long-range-arena code bases and added the normalized attention model and the selective SSM models, respectively. The code for both benchmarks is provided on GitHub for MQAR [^7] https://github.com/IntelligentControlSystems/dsf-mqar and for LRA [^8] https://github.com/jsie7/ssm-benchmark separately.
J.1 MQAR experiments
@@
    Attention: softmax attention [2], linear attention [14], normalized attention (22). For all three attention functions, we use a standard GPT-2 style multi-headed Transformer architecture, where we replace the attention block with the respective attention function. The three attention functions are defined in Section 2.1. For all MQAR runs we use a single attention head.
@@
    State space model: S6 [7], SSD [8]. For both SSM variants, we use a standard GPT-2 style single-headed Transformer architecture, where we replace the attention block with the respective SSM variant. This means for S6 and SSD, we do not implement the pre-convolution on the input or the SiLU activations; but just the S6 and SSD blocks. We do this to ensure a fair comparison of the backbones (sequence mixers) irrespective of the higher-level architecture. The S6 and SSD blocks are defined in Section 2.2 and we use the provided code base [^9] https://github.com/state-spaces/mamba to implement it.
@@
    RNN: qLSTM [9], modified qLSTM. We embed both qLSTM variants in a standard GPT-2 style single-headed Transformer architecture, where we replace the attention block with the qLSTM. We do this to ensure a fair comparison of the backbones (sequence mixers) irrespective of the higher-level architecture. The standard qLSTM is defined in Section 2.3 and the modified qLSTM is the same as the standard qLSTM but with modified state transition (23), i.e., a modified forget gate.
@@
  • Width and depth: For all runs, we use two layers (each with a sequence model and a MLP, interleaved with layer normalization). The model dimensions $d$, state expansion $n$, sequence length $L$, and number of KV pairs are swept according to the experiment (see Section 4). This is the same setup as in [24].
  • Position information: Positional embeddings [60] are used for the attention and RNN architecture classes, but not for the SSM architecture classes. This is the same setup as in [24].
  • Data: Each model is trained on 100,000 datapoints and evaluated on 3,000 datapoints. The data and its order are constant for all runs. This is the same setup as in [24].
@@
resulting in a total of $320$ experiments. We only report the results of the best performing learning rate; the full results of all experiments are stated in Appendix L.
@@
We evaluate the following two architecture classes:
  1. 1.
    Attention: softmax attention [2], linear attention [14], normalized attention (22). For all three attention functions, we use a standard GPT-2 style multi-headed Transformer architecture, where we replace the attention block with the respective attention function. The three attention functions are defined in Section 2.1. To ensure a fair comparison, we keep all hyperparameters of the three attention models constant except the attention function.
  2. 2.
    State space model: S6 [7]. We use a standard GPT-2 style single-headed Transformer architecture, where we replace the attention block with the S6 block. This means, we do not implement the pre-convolution on the input or the SiLU activations; but just the S6 block. We do this to ensure a fair comparison of the backbones (sequence mixers) irrespective of the higher-level architecture. The S6 block is defined in Section 2.2 and we use the provided code base [^11] https://github.com/state-spaces/mamba to implement it.
For all LRA runs, we use the following training protocol:
  • Optimizer and schedule: Linear warmup with duration of 10% and AdamW optimizer [59].
  • Position information: Positional embeddings [60] are used for the attention architecture classes, but not for the SSM architecture classes.
  • Data: Each model is trained on the standard datasets provided with the LRA benchmark.
The exact hyperparameters for each LRA task and each model are reported in the publicly available code base. [^12] https://github.com/jsie7/ssm-benchmark Note that we do not optimize the hyperparameters, i.e., the reported accuracies might be lower than in the original LRA paper [3].
Performed Experiments
We run the three attention models and the S6 models on the LRA tasks ListOps, Text, Retrieval, Image, and Pathfinder-32, which are summarized below; for the full details we refer to [3].
  1. 1.
[…]
  2. 2.
    Text Classification (Text): This task evaluates a model’s ability to capture the tone of long tokenized texts. The dataset consists of IMDb movie reviews, which need to be classified as negative or positive in tone. The task is a binary classification task with maximal input lengths of 4k.
  3. 3.
    Document Retrieval (Retrieval): This task evaluates a model’s ability to compress long sequences into representations that are suitable for similarity matching. The dataset consists of tokenized papers published by the American Academy of Neurology (AAN), which need to be classified in having a citation link or not. The task is a binary classification task with maximal input lengths of 8k.
  4. 4.
    Image Classification (Image): This task evaluates a model’s ability to learn 2D spatial relations from a 1D vector. The dataset consists of vectorized images, which depict one of ten possible classes, e.g. a horse or a car. The task is a ten-way classification task with maximal input lengths of 1k.
  5. 5.
    Long-Range Spacial Dependency (Pathfinder-32): This task evaluates a model’s ability to learn spacial dependencies in a vectorized image. The dataset consists of images, which depict two circles and multiple dashed paths. The goal is to evaluate whether the two circles are connected by any of the present paths or not. The task is therefore a binary classification task with maximal input lengths of 2k.
J.3 WikiText-103 experiments
Training Details
We use the 70M parameter Pythia architecture (Pythia70M) [61]. [^14] https://github.com/EleutherAI/pythia For softmax attention we use the standard Pythia attention block, while for linear attention [14] and normalized attention (22) we replace the attention block in the Pythia architecture with the respective attention functions defined in Sections 2.1 and 4.2.